### v3

# SWE-Sharp-Bench: A Reproducible Benchmark for C# Software Engineering Tasks

Abstract
AI coding agents have shown great progress on Python software engineering benchmarks like SWE-Bench, and for other languages like Java and C in benchmarks like Multi-SWE-Bench. However, C# – a prominent enterprise language ranking #5 in the TIOBE index – remains absent from such benchmarks. We introduce SWE-Sharp-Bench, a reproducible software engineering benchmark for C# featuring 150 instances from 17 repositories. Evaluating identical model-agent configurations across languages reveals a significant performance gap: while 70% of Python tasks in SWE-Bench Verified are solved, only 40% of our C# tasks are resolved. We open-source SWE-Sharp-Bench and our entire curation pipeline.
11footnotetext: Equal Contribution.
I Introduction
Large Language Model-powered automated software engineering has gained substantial attention due to its potential for increasing developer productivity. These models now enable code completion, unit test generation, documentation generation, interactive chat interfaces, and – more recently – autonomous coding agents [1, 2]. This has necessitated increasingly sophisticated benchmarks, moving from simple function-level code completion to evaluating these models on their ability to autonomously solve real-world software engineering problems. SWE-Bench [3] and SWE-Bench Verified [4] have become the most widely-used benchmarks for evaluating latest models and agents on software engineering tasks. Other software engineering benchmarks – Multi-SWE-Bench [5], SWE-Bench Multilingual, and SWE-PolyBench – cover additional programming languages beyond Python.
However, examining the TIOBE Programming Community Index – “an indicator of the popularity of programming languages” [^1] https://www.tiobe.com/tiobe-index/ – reveals systematic gaps in benchmark coverage. The top 8 languages by popularity are: Python (#1, covered by SWE-Bench), C++ (#2, covered by Multi-SWE-Bench), C (#3, covered by Multi-SWE-Bench), Java (#4, covered by Multi-SWE-Bench and SWE-PolyBench), C# (#5, no coverage), JavaScript (#6, covered by Multi-SWE-Bench and SWE-PolyBench), Visual Basic (#7, no coverage), and Go (#8, covered by Multi-SWE-Bench). Notably, the entire .NET ecosystem – including both C# and Visual Basic – remains absent from software engineering benchmarks, despite their high rankings. This gap is particularly striking given C#’s importance to enterprise software development. [^2] https://dotnet.microsoft.com/en-us/platform/customers The absence of .NET languages limits our understanding of how models and coding agents perform with C#’s unique characteristics.
In this paper, we introduce SWE-Sharp-Bench, the first software engineering task benchmark for the C# and .NET ecosystem, comprising 150 curated instances, by adapting SWE-Bench’s methodology. Our curation pipeline tackles .NET-specific challenges including sophisticated dependency management, multi-version compatibility, and cross-platform development to ensure automated creation of reproducible containerized environments. Evaluating these instances on leading models from OpenAI and Anthropic using popular agent frameworks (SWE-Agent and OpenHands [6]), we reveal that C# presents significant challenges for current models, with performance gaps that appear to stem from the relatively high complexity of typical changes in C# projects.
Figure 1: Curation pipeline
II Related Work
Early code-generation benchmarks such as HumanEval [7] and MBPP [8] established the standard for code-generation evaluation. This approach was subsequently scaled and generalized in HumanEval-XL, MBXP, and MultiPL-E [9] (which includes C#). However, these benchmarks still primarily evaluated small, self-contained programming tasks. A shift toward repository-level software-engineering evaluation emerged with SWE-Bench and SWE-Bench Verified, which assess real-world pull requests from open-source repositories. This methodology has since broadened along several axes. In the multilingual direction, Multi-SWE-Bench provides 1,632 instances across seven languages (Java, JavaScript, Go, Rust, C, and C++), SWE-Bench Multilingual [^3] https://kabirk.com/multilingual contributes 300 instances spanning nine languages (C, C++, Java, JavaScript, Go, Rust, TypeScript, PHP, and Ruby), and SWE-PolyBench focuses on JavaScript, TypeScript, Python, and Java. In parallel, GitBug-Java [10] develops a similar repository-level benchmark for Java, using GitHub Actions to ensure reproducible builds. In addition to text-only contexts, SWE-Bench Multimodal [11] extends evaluation to issues that require visual understanding. Beyond curated issues from open-source repositories, SWE-Lancer [12] introduces end-to-end tasks sourced from freelancing platforms. Recent work, such as SWE-Smith [13], also explore scaling instance creation via synthetic data generation with LLMs. Despite these multilingual and methodological advances, C# remains underrepresented in repository-level software-engineering evaluation.
III Building SWE-Sharp-Bench
Figure 2: Distribution of #Files, #Hunks and #Lines in Patches across Languages
SWE-Sharp-Bench consists of 150 issue-resolving tasks carefully curated from 17 popular and actively maintained C# GitHub repositories. The tasks are mapped to GitHub issues which either reports bugs or requests a new feature.
III-A Benchmark Construction
1. Repository Selection: From the top 100 C# GitHub repositories, we retain projects with at least 5,000 stars, active maintenance in the past 6 months, and verified build viability.
2. PR Scraping & Attribute Filtering: We scrape the 1,000 most recent PRs per selected repository and retain only that (1) reference at least one GitHub issue, (2) modify test files and (3) were successfully merged into repository’s default branch.
3. Environment Determination: For each PR, we auto-generate a Dockerfile by parsing .github/workflows, *.sln, *.csproj, global.json and .env to infer NuGet dependencies, build targets and environment variables. The generator reconciles version declarations and handles .NET-specific hurdles (NuGet/MSBuild asset selection, multi-targeting conflicts, deprecated runtimes, etc). We validate by building and running the container.
4. Execution-based Filtering: For each PR, we run tests at (1) base , (2) base + test patch and (3) base + test + fix, retaining only cases with pass $\rightarrow$ fail $\rightarrow$ pass; all others are treated as flaky and omitted.
5. Manual Verification: Each candidate PR was independently annotated and cross-checked by the first two authors with standards similar to SWE-Bench Verified. We flag (1) under specified problem statements and (2) inadequate tests ( overly narrow or misaligned ). Only PRs that pass this review are considered for the final benchmark.
Full details about the benchmark construction process are discussed in Appendix.
III-B Benchmark Characterization
III-B1 Features of SWE-Sharp-Bench
SWE-Sharp-Bench represents a variety of tools and applications. The repositories can be categorized into Data & Storage (4), API Infrastructure (3), User Interface(3), Development Tools (5) and Multimedia Processing (2). The categorization is done by manual inspection of the repository descriptions (see Appendix Table II). The instances are categorized into three primary categories Bug-Fixes (91), Feature Requests (47) and Others (12). The categorization process is discussed in Appendix A-C . 53% of the instances are created in 2024 and 90% of the instances are created after 2023.
III-B2 Characteristics of SWE-Sharp-Bench versus Other Benchmarks
Language and Benchmark Selection: Before attributing performance differences to model or agent limitations, we need to understand the difference in characteristics of the current benchmarks. We select Python, the essential baseline with highest representation in research interests and Java which represents a natural comparison point for C# as it shares similar properties like static typing, complex build systems and dependency management. We use SWE-Bench-Verified for Python and Multi-SWE-Bench for Java.
Patch Complexity Analysis: We adopt the patch complexity metrics introduced in Multi-SWE-Bench, which measures static properties across three dimensions:
$\bullet$ Patch-level metrics: Files modified (change breadth), hunks per patch (modification granularity), lines added/removed (change magnitude).
$\bullet$ Repository metrics: Total files and lines of code.
$\bullet$ Task specification: Token length of the problem statement.
We extract these metrics for the 150 SWE-Sharp-Bench instances and 500 SWE-Bench Verified instances. For Java, we use the metrics reported in the Multi-SWE-Bench paper. Figure 2 demonstrates the distribution of files modified, number of hunks and number of lines added in patches across languages. In the Appendix, Tables III, IV and V summarize repository-level statistics. Our analysis using these metrics reveal:
Repository Scale: C# projects range from roughly 5k LoC to 1.47 M LoC (median $\approx 235\,\text{}k$), whereas the largest Python repository in SWE‑Bench Verified is 383 k LoC and the largest Java repository in Multi‑SWE‑Bench reaches 443 k LoC. Bigger repositories translate to a broader surface area for the agent to search through.
Change Locality: Python fixes are usually surgical (mean=1.24, median=1 file); Java shows moderate spread (mean=2.96, median=2); and C# has the broadest distribution (mean=4.88, median=2), combining many small fixes with a long tail of multi-file changes, creating a diverse mix.
Modification Granularity: Patch depth increases progressively from Python (2.4 hunks and 14.3 lines on average ) to Java (6.26/89.27) to C# (10.0/131.1), as shown in Figure 2, middle. C# exhibits the most diverse distribution: like Python and Java, many patches are small, but C# also includes substantially larger modifications. Manual inspection of high patch-count C# instances revealed these typically involve refactoring operations and coordinated multi-file edits.
Task Specification: Problem‑statement length varies markedly by language. Java issues can exceed $\sim$1.6 k tokens, Python issues generally stay below 450, and C# issues are often under 150. Patches with short descriptions may be difficult for agents to solve if they are insufficiently detailed.
Takeaways: C# exhibits the most complex static patch properties among the three languages, and SWE-Sharp-Bench provides a diverse mix of task complexities.
IV Experiments and Results
TABLE I: Resolution Rates (%) of Open-Source agents across various models on SWE-Sharp-Bench
Model SWE-Agent OpenHands
GPT-4o 11.3 8.0
GPT-4.1 22.0 23.3
GPT-5 43.3 47.3
o3-mini 19.3 19.3
o4-mini 25.3 26.0
o3 33.3 35.0
Claude Sonnet 3.5 20.0 22.6
Claude Sonnet 3.7 31.3 31.3
Claude Sonnet 4 44.7 40.6
Agents & Models: We evaluate two popular agent systems: SWE-Agent and OpenHands. We test each system across multiple leading language models from OpenAI and Anthropic. Since these frameworks were originally designed for Python repositories, we adapt their prompts for C# projects. Each agent receives a single attempt per instance with a 2-hour timeout limit. due to budget constraints, we conduct a single attempt per instance, retrying only when infrastructure failures (e.g., rate limits, API errors ) occur to ensure at least one valid attempt per instance.
Evaluation Metrics: We use Resolution rate as the primary metric, the percentage of instances successfully resolved by each agent. An instance is considered resolved when the agent’s generated patch passes all the required tests.
Results: Table I summarizes performance on SWE-Sharp-Bench of different OpenAI and Anthropic models using SWE-Agent and OpenHands. Across all configurations, OpenHands + GPT-5 performs the best with 47.3 % resolution rate. Table VI in Appendix, demonstrates performance on SWE-Bench Verified and Multi-SWE-Bench’ Java subset on identical model-agent combinations. More results are discussed in the Appendix.
Figure 3: Logistic regression coefficients and their influence on resolution rate (baseline: C#, GPT-4o). Significance: *** p < 0.001, ** p < 0.01, * p < 0.05.
IV-A Which factors affect agent performance ?
We observe performance gaps between Python when compared to C# and Java for identical model-agent configurations, e.g. on SWE-Agent + Claude Sonnet 3.7 , Resolution rate for Python is 62.40 % , 30.67 % for C# and 14.68 % for Java. Refer Table 5 in the appendix for more details. The likelihood of successful instance resolution depends on multiple factors: how spread out the patch is ( hunks and files in the patch), how big the patch is ( patch lines), which model was used, or which agent was used. We observe that agents perform significantly better on Python compared to C# and Java, but it’s unclear how much of this performance gap is correlated with the instance’s static properties versus the choice of programming language itself. To understand which factors most strongly influence resolution success, we use logistic regression analysis. We conduct this analysis using instance-level data from SWE-Agent runs with GPT-4o, Claude Sonnet 3.5 and Claude Sonnet 3.7 models, the combinations with complete instance level resolution data available across all three benchmarks.
Figure 3 shows the regression coefficients with GPT-4o and C# as baseline categories. As expected, model choice has the strongest influence, with newer models significantly outperforming GPT-4o. More complexity decreases success: patch hunks, lines, and files negatively affect performance, indicating that widespread edits across multiple locations are more challenging for agents to resolve. Programming language shows a substantial effect, with Python significantly easier than the C# baseline. This analysis confirms that Python is the easiest and – after controlling for complexity and model – Java and C# are similarly difficult.
V Limitations
While SWE-Sharp-Bench provides a diverse mix of tasks, it has few limitations. First, with 150 instances, it is smaller than SWE-Bench Verified’s 500, though comparable to individual language subsets in Multi-SWE-Bench. Second, unlike SWE-Bench Verified and Multi-SWE-Bench, we do not provide manual annotations for difficulty, though patch complexity metrics offer objective complexity indicators. Finally, as with any benchmark derived from public data, the included data might have been used for any recent LLM training.
VI Benchmark & Data Availability
The Appendix contains detailed analysis, a breakdown of the benchmark, and a deep dive into the performance results. Benchmark data is available at HuggingFace444huggingface.co/datasets/microsoft/SWE-Sharp-Bench. The curation pipeline code and agent trajectories are available at GitHub555github.com/microsoft/prose/tree/main/misc/SWE-Sharp-Bench.
Appendix A Appendix
A-A C# Background
For readers unfamiliar with C#, understanding its project structure is essential to appreciating the unique challenges it presents for automated environment building. While Python and Java projects can certainly be complex with their own packaging systems, C# introduces additional layers of hierarchy through its project system, A .sln solution file acts as a master container, similar to a workspace, that can contain multiple .csproj (C# project) files, each defining a separate component like a main application, test suite or shared library. This multi-layered structure becomes particularly challenging with C#’s multi-targeting feature, where a single project can be compiled for different .NET versions simultaneously (eg, .NET 6.0, .NET Framework 4.8 and .NET standard 2.0). Imagine if a Python project needed to maintain compatibility with Python2.7, 3.8 and 3.11 simultaneously within the same code base, with different APIs for each version, this is routine in C#. This architectural complexity, combined with various build configurations, platform targets and the tight coupling between Visual Studio tooling and project files, make C# repositories more challenging to automatically analyze and test compared to the relatively flat structure of Python packages or Java’s more uniform build systems like Maven or Gradle. These intricate inter-dependencies mean that what might be a simple pip install and pytest in Python becomes a complex orchestration of MSBuild targets, NuGet package restoration and framework-specific test runners in C#.
A-B Benchmark Construction
1. Repository Selection: We select a high-quality set of GitHub repositories through a multi-stage filtering process.
Popularity Filtering: We identified the top 100 C# repositories on GitHub ranked by GitHub stars. From this initial set, we applied a minimum threshold of 5,000 stars.
Active Maintenance: We verify that selected repositories demonstrate active development and maintenance within the last 6 months by manually analyzing three indicators: commit history frequency, merged pull requests, and issue creation activity.
Build Viability: Verification by either minimal manual setup or executing local GitHub Actions workflows with repo’s latest commit. Only repositories that successfully build and pass their test suites at the latest commit were considered.
2. Pull Request Scraping and Attribute-based Filtering: We start by scraping the 1000 most recent PRs from each selected repository. All PRs are then filtered by the following criteria:
Linked with at least one GitHub issue: The PR must reference at least one GitHub issue to ensure it addresses either a bug report or a feature request.
Changes to test files: The PR must include modifications to test files, indicating the author contributed to testing to ensure the issue is resolved. We identify test files using keyword pattern matching for "test" or "testing" in filenames or filepaths, which captures the majority of test files in our dataset.
Merged in main branch: The PR must be successfully merged into the main branch, indicating the PR was thoroughly reviewed and approved by repository maintainers.
3. Environment Determination: To ensure consistently reproducible environments, for each PR we automatically construct a Dockerfile which parses .github/workflows, .sln, .env, global.json and .csproj files to automatically determine the project’s environment variables and dependencies. This automated construction addresses several unique challenges in .NET containerization: managing complex .NET dependency resolution compared to simpler package managers like pip, handling multiple .NET framework versions within single containers, dealing with deprecated .NET versions that require specific base images, and resolving projects that target multiple .NET versions simultaneously. Our system automatically detects environment variables from various configuration sources and reconciles version conflicts across different project files.
We validate each Dockerfile by creating the Docker image and launching the container to ensure consistent environment setup. Any failed builds are manually analyzed for missing dependencies, misconfigurations, or version conflicts. Such issues are fixed if they require minimal modifications to the Dockerfile that do not involve significant human effort. Otherwise, such PRs are discarded from our dataset.
4. Execution-based Filtering: For each PR, we run tests in three states: at the base commit, after applying the test patch at the base commit and finally applying both fix and test patch at the base commit. We filter PRs to include only those demonstrate this pattern: all test pass at the base commit, at least one test fails after applying the test patch and again all tests pass when we apply both fix and test patch. Any other scenarios are marked flaky and discarded from the dataset.
Implementing this execution-based validation for .NET projects required addressing different challenges compared to ecosystems like Python. Unlike Python’s widely used and more standardized pytest ecosystem, different .NET projects utilize different testing frameworks (NUnit, XUnit, MSTest) each with different execution patterns and output formats. Our system automatically detects and handles these testing frameworks, unifies their logging mechanisms and normalizes their output formats to support a common evaluation harness across all repos.
5. Manual Verification: Finally, each PR is annotated and cross-verified by first two authors. The annotations were carried out with standards set similar to SWE-Bench Verified. The candidate PRs are checked for:
Underspecified Problem Statements: We annotate if the problem statement is underspecified, leading to ambiguity on what the problem is or how it should be solved.
Thorough Test Cases: We check whether the test cases introduced in the PR are overly specific, narrow, or unrelated to the actual problem being solved.
Instances which pass this manual inspection were selected for the final benchmark.
Figure 4: SWE-Sharp-Bench Curation Process
TABLE II: Detailed Repository Descriptions
Repository Description
App-vNext/Polly Polly is a .NET resilience and transient-fault-handling library that allows developers to express policies such as Retry, Circuit Breaker, Timeout, Bulkhead Isolation, and Fallback in a fluent and thread-safe manner.
AvaloniaUI/Avalonia Develop Desktop, Embedded, Mobile and WebAssembly apps with C# and XAML. The most popular .NET UI client technology
JoshClose/CsvHelper Library to help reading and writing CSV files
MessagePack-CSharp/MessagePack-CSharp Extremely Fast MessagePack Serializer for C#(.NET, .NET Core, Unity, Xamarin). / msgpack.org[C#]
SixLabors/ImageSharp A modern, cross-platform, 2D Graphics library for .NET
StackExchange/
StackExchange.Redis
General purpose redis client
ThreeMammals/Ocelot .NET API Gateway
ardalis/CleanArchitecture Clean Architecture Solution Template: A proven Clean Architecture Template for ASP.NET Core 9
autofac/Autofac An addictive .NET IoC container
devlooped/moq The most popular and friendly mocking framework for .NET
dotnet/BenchmarkDotNet Powerful .NET library for benchmarking
dotnet/efcore EF Core is a modern object-database mapper for .NET. It supports LINQ queries, change tracking, updates, and schema migrations.
gui-cs/Terminal.Gui Cross Platform Terminal UI toolkit for .NET
jellyfin/jellyfin The Free Software Media System - Server Backend & API
restsharp/RestSharp Simple REST and HTTP API Client for .NET
serilog/serilog Simple .NET logging with fully-structured events
spectreconsole/spectre.console A .NET library that makes it easier to create beautiful console applications.
A-C Benchmark Characterization
Issue Type Categorization and Distribution
We systematically categorize instances into three primary types: Bug-Fix (97), Feature Requests (41), and Refactors (12), with further subdivision into secondary subtypes. Bug-Fix issues have Logic (54), UI (28), Logging (6), Concurrency (2), Compatibility (2), Stoage (2), API (1), Networking (1) and Security (1). Feature Requests have API (16), Logic (6), UI (6), Configuration (4), Logging (4), Storage (2), Documentation (1), Performance (1) and Compatibility (1) as sub-types. For this categorization, we selected the labels by manual inspection of few sampled tasks. After deciding on the labels, we used GPT-5 to provide a label using problem statement, patch and the repo’s description. 5
TASK: Classify software development issues into categories based on the provided information. INPUT PROVIDED: 1. Repository Information: Details about the software project 2. Problem Statement: Description of the issue or feature request 3. Code Changes (Patch): The actual code modifications made to address the issue CLASSIFICATION CATEGORIES: PRIMARY LABEL (required - select exactly one): -> Feature: Adding new functionality, enhancements, or feature requests -> Bug: Fixing incorrect behavior, errors, or defects -> Other: Documentation updates, build system changes, infrastructure, etc. SECONDARY LABEL (required - select the most relevant one): - UI: User interface, visual elements, controls, rendering issues - Performance: Speed optimization, resource usage improvements - Security: Authentication, authorization, vulnerability fixes - Storage: Database operations, file system, data persistence - Logging: Logging systems, telemetry, monitoring, tracing - API: External interfaces, endpoints, API integration - Configuration: Settings, options, setup, environment configuration - Testing: Test frameworks, test utilities, validation logic - Documentation: Documentation updates, comments, help text - Build: Compilation, packaging, deployment, build system - Compatibility: Version compatibility, platform support - Logic: Business logic, algorithms, calculations, core functionality - Memory: Memory management, memory leaks, allocation issues - Concurrency: Threading, async operations, race conditions - Networking: HTTP, TCP, network protocols, connectivity - Other: If none of the above categories fit well INSTRUCTIONS: - Analyze the repository context to understand the project domain - Read the problem statement to understand what needs to be addressed - Examine the code changes to see what was actually implemented/fixed - Choose the PRIMARY label based on whether this adds functionality (Feature), fixes a problem (Bug), or is maintenance work (Other) - Choose the SECONDARY label based on the technical area most affected by the changes OUTPUT FORMAT (JSON only): {{ "primary": "Feature|Bug|Other", "secondary": "UI|Performance|Security|Storage|Logging|API|Configuration|Testing|Documentation|Build|Compatibility|Logic|Memory|Concurrency|Networking|Other" }} --- {repo_description} Problem Statement: {problem_statement} Code Changes (Patch): {patch} Analyze the above information and provide your classification in the exact JSON format specified.
Figure 5: Prompt Template used for Issue Type Categorization
A-D Instance Statistics
TABLE III: Statistics of SWE-Sharp-Bench
Org/Repo Repository Instance Gold Patch Test Patch Unit tests
#Files #LoC #Count Avg. #Tokens Avg. #Lines Avg. #Hunks Avg. #Files Avg. #Lines Avg. #Hunks Avg. #Files #Avg. F2P #Avg. P2P
App-vNext/Polly 790 84.8k 14 176.64 98.57 13.64 7.50 123.79 14.21 6.21 1.64 5.36
AvaloniaUI/Avalonia 3532 436.5k 41 287.02 147.95 8.80 3.73 63.51 3.54 1.32 2.29 4.10
JoshClose/CsvHelper 461 44.8k 1 192.00 369.00 35.00 22.00 423.00 36.00 30.00 13.00 10.00
MessagePack-CSharp/MessagePack-CSharp 688 80.8k 18 292.72 41.61 7.11 2.28 117.78 3.39 2.33 1.67 0.56
SixLabors/ImageSharp 1997 248.4k 8 256.38 56.00 4.62 2.62 42.75 4.62 2.38 0.75 6.88
StackExchange/StackExchange.Redis 341 72.8k 3 406.67 50.67 9.67 5.00 104.33 15.33 4.67 2.67 0.00
ThreeMammals/Ocelot 735 53.5k 4 146.50 498.75 44.00 24.50 699.50 31.50 17.25 10.25 7.50
ardalis/CleanArchitecture 256 5.0k 3 125.33 34.33 4.33 4.33 45.00 6.67 5.67 3.33 0.33
autofac/Autofac 577 45.8k 6 411.83 134.83 10.67 5.17 83.33 2.00 1.33 3.83 13.33
devlooped/moq 242 40.8k 4 287.75 34.25 4.50 3.00 55.50 2.00 1.00 3.00 0.00
dotnet/BenchmarkDotNet 955 67.0k 8 291.88 54.75 6.00 3.38 71.88 2.12 1.12 3.62 14.00
dotnet/efcore 5468 1479.1k 9 232.78 565.00 20.11 8.22 174.11 11.22 5.89 1.00 14.22
gui-cs/Terminal.Gui 938 221.7k 1 113.00 22.00 3.00 2.00 5.00 1.00 1.00 1.00 0.00
jellyfin/jellyfin 1943 224.7k 3 679.00 23.00 2.00 2.00 30.33 1.33 1.00 1.00 6.67
restsharp/RestSharp 226 15.9k 5 192.00 66.60 11.60 5.20 37.40 4.40 2.40 1.80 2.40
serilog/serilog 214 21.2k 12 257.08 80.50 7.42 3.25 84.25 8.08 4.92 3.25 4.25
spectreconsole/spectre.console 729 62.5k 10 172.10 55.60 7.40 4.70 68.20 9.10 6.40 1.20 0.10
TABLE IV: Statistics of SWE-Bench Verified
Org/Repo Repository Instance Gold Patch Test Patch Unit tests
#Files #LoC #Count Avg. #Tokens Avg. #Lines Avg. #Hunks Avg. #Files Avg. #Lines Avg. #Hunks Avg. #Files #Avg. F2P # Avg. P2P
astropy/astropy 526 160.9k 22 402.91 27.77 2.27 1.23 38.27 2.36 1.18 2.23 167.55
django/django 833 114.1k 231 196.99 11.81 2.06 1.2 24.98 2.39 1.42 4.58 83.32
matplotlib/matplotlib 265 124.1k 34 331.76 9.26 2.09 1.18 19.5 1.53 1.03 1.82 378.29
mwaskom/seaborn 63 23.4k 2 237.0 13.5 3.0 1.5 18.5 1.5 1.5 2.0 171.0
pallets/flask 25 6.8k 1 45.0 3.0 1.0 1.0 5.0 1.0 1.0 1.0 59.0
psf/requests 21 4.4k 8 224.88 3.62 1.5 1.0 6.0 1.0 1.0 3.88 100.25
pydata/xarray 132 82.0k 22 392.68 18.0 2.36 1.23 22.68 1.55 1.18 2.18 657.73
pylint-dev/pylint 1082 43.5k 10 450.8 24.7 3.7 2.1 33.8 1.9 1.2 4.2 46.3
pytest-dev/pytest 85 28.5k 19 336.42 22.58 2.42 1.11 48.32 3.11 1.26 1.84 64.74
scikit-learn/scikit-learn 423 179.7k 32 379.81 12.31 2.97 1.06 20.22 1.72 1.06 1.22 64.19
sphinx-doc/sphinx 226 66.8k 44 249.82 16.84 2.45 1.25 25.52 2.34 1.55 1.2 23.36
sympy/sympy 858 383.4k 75 162.27 16.63 3.56 1.44 14.63 2.11 1.28 1.25 51.99
TABLE V: Statistics of Multi-SWE-Bench (Java only)
Repository Instance Fix patches Unit tests
Org/Repo #Files #LoC #Num Avg. #Tokens Avg. #Lines Avg. #Hunks Avg. #Files #A2P2P #A2F2P #A2N2P
Java
alibaba/fastjson2 4244 443.8k 6 459.2 10.5 1.3 1.2 1243.5 0.8 1020.5
elastic/logstash 562 59.9k 38 1600.4 212.3 10.0 4.6 554.7 1.9 256.2
mockito/mockito 986 84.0k 6 315.2 92.5 10.3 4.7 97.2 1.0 3.8
apache/dubbo 3939 402.1k 3 774.0 9.3 3.0 1.3 2.0 57.0 0.0
fasterxml/j-core 366 105.7k 18 304.7 33.8 4.8 2.1 2.0 85.6 0.0
fasterxml/j-dbind 1230 217.5k 42 621.5 35.1 3.9 2.1 2.0 73.8 0.0
fasterxml/j-dfmt-xml 206 23.0k 5 1071.8 98.4 10.4 3.2 2.0 94.2 0.0
google/gson 261 48.0k 5 365.8 35.8 4.6 1.8 2.0 62.6 0.0
google-ct/jib 604 75.5k 5 1094.6 15.2 3.2 2.6 2.0 96.2 0.0
TypeScript
darkreader/darkreader 189 26.2k 2 749.5 13.0 2.0 1.5 41.0 3.5 0.0
mui/material-ui 27632 698.6k 174 508.6 331.2 20.2 12.0 5001.3 2.3 836.8
type/core 509 128.2k 48 694.8 22.9 3.5 1.9 2920.4 3.0 0.0
JavaScript
ag/gh-rtdme-stats 69 11.8k 19 287.1 123.6 13.5 4.8 108.9 3.5 3.4
axios/axios 166 21.0k 4 490.8 179.5 7.8 4.0 68.5 1.2 0.0
expressjs/express 142 17.3k 4 177.5 7.2 2.2 1.5 808.2 1.5 65.2
iamkun/dayjs 324 17.1k 56 325.6 21.7 2.7 2.0 60.4 1.2 3.2
Kong/Insomnia 526 182.0k 1 709.0 1.0 1.0 1.0 105.0 1.0 0.0
sveltejs/svelte 2800 105.9k 272 618.9 72.0 8.4 4.0 4904.2 5.5 0.0
Go
cli/cli 737 165.1k 397 347.6 103.8 9.0 3.9 1997.0 2.9 31.0
grpc/grpc-go 981 260.8k 16 276.1 81.8 7.7 2.8 230.4 0.6 6.6
zeromicro/go-zero 960 117.6k 15 205.2 52.4 4.9 2.7 1318.9 0.3 43.9
Rust
BurntSushi/ripgrep 98 45.4k 14 553.7 1604.9 21.9 7.5 233.2 1.1 8.1
clap-rs/clap 321 70.4k 132 987.0 147.1 15.7 4.7 489.5 3.1 378.8
nushell/nushell 1479 264.2k 14 795.6 155.0 10.6 4.3 798.6 2.6 336.6
rayon-rs/rayon 191 36.9k 2 153.5 637.5 5.5 2.0 113.5 0.5 171.0
serde-rs/serde 188 36.5k 2 171.5 72.5 3.0 3.0 0.0 0.0 294.5
sharkdp/bat 83 22.0k 10 638.2 239.5 14.1 5.9 152.7 1.7 33.6
sharkdp/fd 24 6.7k 14 167.8 55.8 7.8 4.5 186.5 1.1 0.0
tokio-rs/bytes 33 11.9k 5 188.0 45.0 5.6 1.8 23.2 0.4 91.6
tokio-rs/tokio 727 141.5k 25 590.0 139.8 10.6 3.5 26.6 0.0 287.4
tokio-rs/tracing 241 60.9k 21 472.0 597.2 39.3 7.1 30.8 0.2 182.0
C
facebook/zstd 276 119.8k 29 496.6 67.6 10.9 3.0 0.8 0.5 5.6
jqlang/jq 80 43.0k 17 429.8 26.1 2.7 1.8 27.2 1.0 0.1
ponylang/ponyc 285 80.2k 82 480.2 205.4 15.6 5.7 997.6 1.9 388.8
C++
catchorg/Catch2 399 58.0k 12 357.3 469.0 15.4 8.2 19.9 0.7 17.6
fmtlib/fmt 25 36.4k 41 397.7 36.8 3.0 1.1 9.3 0.0 9.3
nlohmann/json 477 124.7k 55 905.5 405.8 27.9 6.5 26.5 0.0 42.9
simdutf/simdutf 455 229.7k 20 320.2 768.5 35.5 11.0 18.6 0.0 41.5
yhirose/cpp-httplib 33 50.9k 1 240.0 1.0 1.0 1.0 272.0 1.0 0.0
A-E Extended Results
We use this section to provide some additional results which include performance of different models with respect to different dimensions. All the SWE-Sharp-Bench agent runs were scheduled by us, using the prompt template mentioned in Figure 6. Table VI provides resolution rate for model + agent configurations combinations across all three benchmarks. Resolution rates reported for SWE-Bench Verified and Multi-SWE-Bench are obtained from their respective public leader-boards  [^6] https://www.swebench.com/, https://multi-swe-bench.github.io/. SWE-Agent + Claude 4 Sonnet is one exception to this, this was scheduled by us. Due to resource constraints we were only able to do a single entire benchmark with this configuration.
You are a helpful assistant that can interact with a computer to solve tasks. <uploaded_files> {{working_dir}} </uploaded_files> Ive uploaded a C# repository in the directory {{working_dir}}. Consider the following PR description: <pr_description> {{problem_statement}} </pr_description> Can you help me implement the necessary changes to the repository so that the requirements specified in the <pr_description> are met? Ive already taken care of all changes to any of the test files described in the <pr_description>. This means you DONT have to modify the testing logic or any of the tests in any way! Your task is to make the minimal changes to non-tests files in the {{working_dir}} directory to ensure the <pr_description> is satisfied. Follow these steps to resolve the issue: 1. As a first step, it might be a good idea to find and read code relevant to the <pr_description> 2. Create a script to reproduce the error and execute it using the bash tool, to confirm the error 3. Edit the sourcecode of the repo to resolve the issue 4. Rerun your reproduce script and confirm that the error is fixed! 5. Think about edgecases and make sure your fix handles them as well Your thinking should be thorough and so it’s fine if it’s very long.
Figure 6: Prompt Template used for SWE-Agent and OpenHands run on SWE-Sharp-Bench
TABLE VI: AI Agent Benchmark Resolution Rates
Agent Model Resolution Rate (%)
Python - SWE-Bench Verified Java - Multi-SWE-Bench C# - SWE-Sharp-Bench
SWE-agent GPT-4o 26.00 5.11 11.30
Claude Sonnet 3.5 33.60 10.20 20.00
Claude Sonnet 3.7 62.40 14.68 31.30
Claude Sonnet 4 66.60 18.75* 44.70
OpenHands GPT-4o 25.75 5.96 8.00
GPT-4.1 48.60 10.11 23.30
OpenAI o3-mini 43.70 6.29 19.33
OpenAI o3 59.00 21.00 35.00
Claude Sonnet 3.5 53.00 12.73 22.60
Claude Sonnet 3.7 60.60 16.01 31.30
A-E1 Performance by Repository:
In Table VII we provide a breakdown of repository-level performance across 2 dimensions: Agent and Model. We used GPT-4o and GPT-5 from OpenAI, and Claude Sonnet 3.5 and Claude Sonnet 4 from Anthropic. We use these combinations to show progress between model generations.
TABLE VII: Repository-wise Resolve Rate Comparison
OpenHands SWE-Agent Repository GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 dotnet/efcore 0.0 55.6 33.3 44.4 55.6 55.6 44.4 55.6 serilog/serilog 33.3 50.0 33.3 58.3 16.7 41.7 33.3 58.3 spectreconsole/spectre.console 10.0 60.0 40.0 60.0 10.0 60.0 20.0 60.0 SixLabors/ImageSharp 0.0 62.5 25.0 75.0 12.5 62.5 25.0 50.0 autofac/Autofac 16.7 50.0 33.3 50.0 0.0 50.0 33.3 50.0 restsharp/RestSharp 0.0 60.0 20.0 60.0 0.0 60.0 20.0 60.0 dotnet/BenchmarkDotNet 12.5 62.5 25.0 50.0 0.0 62.5 12.5 50.0 devlooped/moq 0.0 50.0 25.0 75.0 0.0 50.0 0.0 75.0 AvaloniaUI/Avalonia 9.8 41.5 24.4 31.7 14.6 36.6 22.0 41.5 StackExchange/StackExchange.Redis 0.0 33.3 0.0 33.3 0.0 33.3 33.3 66.7 App-vNext/Polly 0.0 42.9 21.4 28.6 7.1 42.9 14.3 35.7 MessagePack-CSharp/MessagePack-CSharp 0.0 50.0 5.6 27.8 5.6 44.4 5.6 38.9 ardalis/CleanArchitecture 33.3 33.3 33.3 0.0 0.0 33.3 33.3 0.0 ThreeMammals/Ocelot 0.0 25.0 0.0 50.0 0.0 0.0 0.0 0.0 jellyfin/jellyfin 0.0 33.3 0.0 0.0 0.0 0.0 0.0 33.3 gui-cs/Terminal.Gui 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 JoshClose/CsvHelper 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 OVERALL 8.0 47.3 22.6 40.6 11.3 43.3 20.0 44.7
A-E2 Performance by Issue Creation Year:
Table VIII provides a temporal breakdown that shows resolution for instances across different years. We use the same model + agent configurations mentioned in A-E1.
TABLE VIII: Year-wise Resolve Rate Comparison
OpenHands SWE-Agent
Year GPT-4o GPT-5 Sonnet-3.5 Sonnet-4 GPT-4o GPT-5 Sonnet-3.5 Sonnet-4
2020 0.0 60.0 20.0 80.0 0.0 60.0 0.0 80.0
2021 0.0 100.0 100.0 100.0 0.0 100.0 100.0 100.0
2022 0.0 44.4 22.2 55.6 33.3 44.4 22.2 33.3
2023 7.4 42.6 20.4 40.7 3.7 38.9 18.5 38.9
2024 10.4 50.6 23.4 36.4 15.6 45.5 22.1 48.1
2025 0.0 25.0 25.0 25.0 0.0 25.0 0.0 25.0
TOTAL 8.0 47.3 22.6 40.6 11.3 43.3 20.0 44.7
A-E3 Trajectory Analysis:
In this section we provide trajectory analysis of the same model + agent combinations from the previous sections. Figure 7 visualizes the distribution of the number of turns that were resolved. Figure 8 demonstrates the percentage of instances that were successfully localized by the agent and resolved finally. The criteria of successful localization is the agent lands up at least one of the files from the golden fix patch.
(a) OpenHands + GPT-4o
(b) SWE-Agent + GPT-4o
(c) OpenHands + GPT-5
(d) SWE-Agent + GPT-5
(e) OpenHands + Sonnet 3.5
(f) SWE-Agent + Sonnet 3.5
(g) OpenHands + Sonnet 4
(h) SWE-Agent + Sonnet 4
Figure 7: Comparison of OpenHands vs SWE-Agent performance across different language models on SWE-Sharp-Bench
(a) OpenHands + GPT-4o
(b) SWE-Agent + GPT-4o
(c) OpenHands + GPT-5
(d) SWE-Agent + GPT-5
(e) OpenHands + Sonnet 3.5
(f) SWE-Agent + Sonnet 3.5
(g) OpenHands + Sonnet 4
(h) SWE-Agent + Sonnet 4
Figure 8: Localization to Resolution flow